OpenStudy (anonymous):

Is a triangle with sides of length 6 ft., 21 ft., 23 ft. a right triangle? can someone help me with this one?

OpenStudy (anonymous):

use pythagoras's theorem to verify .

OpenStudy (anonymous):

i got you bro hang on

OpenStudy (anonymous):

a2*b2=c2 right?

OpenStudy (anonymous):

A right triangle obeys the Pythagorean Theorem a^2 + b^2 = c^2. Plug in the values of the three sides and see if it works. c would be the longest side.

OpenStudy (anonymous):

correct as RLTW says

OpenStudy (anonymous):

36*441=529 like this?

OpenStudy (anonymous):

6^2=36 21^2=441 23^2=529 36+441 does not equal 529 so it is not a right triangle

OpenStudy (anonymous):

ohhh ok I see!! Thanks man!!! and thanks Surjithayer!

OpenStudy (anonymous):

\[36+441=507\neq 529\]

OpenStudy (anonymous):

your well come bro